6. Torque Manufacturing forecasts that its production will require 600,000 tons of bauxite over its planning period (365 days). Demand for Torque's products is stable over time. Ordering costs amount to an average of $15 per order. Holding costs are estimated at $1.25 per ton of bauxite. The EOQ for Torque is _____ tons.

7. Following Question 6, what is the total annual cost of inventory?

8. Following Question 6, if it takes 7 days to receive an order from suppliers, at what inventory level should anther order be replaced?
Answer #1

ques 1:

we find EOQ = ((2*ANNUAL DEMAND*ORDER COST)/HOLDING COTS)^1/2

SO WE HAVE DEMAND=600000 TONS

ORDER COST =$15 PER ORDER

HOLDING COST = $1.25 PER TON

USING THE FORMULA WE GET

EOQ UNITS = 3795 UNITS

QUES 2

COST OF INVENTORY = 3795 *1.25 = $4743

QUES 3

REODER QUANTITY

DAILY USAGE = 600000/365 DAYS

REORDER QUANTITY = 600000/365 DAYS * 7 DAYS =11506.85